The particle is thrown up with the velocity of 40m/s. The maximum height reached by particle is?

Grade:9

1 Answers

salmankhan
6 years ago
Rmax =U2/g
=40*40/10
=1600/10
=160 m.
,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,